LASER SPECTROSCOPY OF C_{2}H: A BAND NEAR 6340 CM^{−1}

\underline{DENNIS W. TOKARYK}, Physics Department, University of New Brunswick, Fredericton, NB, Canada E3B 5A3.

C_{2}H is a linear radical found in a variety of terrestrial environments (flames, hydrocarbon photofragments), and is abundant in the interstellar medium. Analysis of its infrared spectrum is complicated by a low-lying electronic state Ã^{2}Π, about 3700 cm^{−1} above the X̃^{2}Σ^{+} ground state. Strong interactions between Ã^{2}Π and X̃^{2}Σ^{+} make it difficult to assess the vibrational character of the upper states in observed spectra. In the course of near-infrared laser spectroscopic studies of a methane-helium discharge, we observed a ^{2}Σ −^{2}Σ band with an origin near 6340 cm^{−1}. Combination difference comparisons confirmed that lower state was the 000 vibrational level of the ground state of C_{2}H. The upper state may be a vibrational state of X̃ ^{2}Σ^{+}, or of Ã^{2}Π if a bending vibration is involved. We will present the rotational analysis of this band, and will discuss possible vibrational assignments.